# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v2

@@ 3 Proposed Approach
To address the issues presented in Section 2.4, we propose an approach that takes advantage of the domain and class observables to constrain the approximate posterior distribution. Our approach ensures that each subset of the latent variables only encodes information that is relevant to the corresponding modality. This constraint in turn limits the amount of information available to the data-driven components of the decoder, preventing them from correcting every discrepancy between the physics-based model and measurements, and from overriding the known physics. This is achieved by imposing a latent bottleneck structure to the model, combined with an adversarial training objective. A detailed description of the model architecture and derivation of the training objective are provided in Section 3.1, followed by a brief discussion in Section 3.2.
@@ 3.1 Detailed description of the model
To minimize the reconstruction error, the encoder tends to maximize the information in the posterior distribution over $\bm{z}$ that can be used to predict $\bm{x}$, $\bm{c}$ and $\bm{y}$, subject to the regularization imposed by the prior distribution. For $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, this information includes features from the input signal $\bm{x}$ that are predictive of $\bm{c}$ and $\bm{y}$, but also irrelevant features that are only predictive of $\bm{x}$. These features can include systematic errors stemming from partial knowledge of the physics, and the influence of unknown confounding factors in the measurements. This superfluous information [Federici2020], i.e. information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that is not predictive of $\bm{c}$ and $\bm{y}$, can enable the data-driven component of the decoder to override the known physics. Motivated by this observation we aim to simultaneously maximize the information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that is predictive of $\bm{c}$ and $\bm{y}$, while minimizing the information that is predictive of $\bm{x}$. This trade-off can be formalized in terms of the Mutual Information (MI), a measure of the dependence between two random variables [Cover2006]. Denoting the MI between $\bm{x}$ and $\bm{z}$ for an encoder parametrized by $\bm{\phi}$ as $I_{\mathrm{\phi}}(\bm{x};\bm{z})$, and introducing the trade-off parameters $\lambda_{\mathrm{c}},\lambda_{\mathrm{y}}$, we define the following relaxed Lagrangian objectives:
$\begin{split}\mathcal{L}_{\mathrm{y}}(\bm{\phi};\lambda_{\mathrm{y}})&=\ I_{\mathrm{\phi}}(\bm{y};\bm{z}_{\mathrm{y}})-\lambda_{\mathrm{y}}I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{y}})\\ \mathcal{L}_{\mathrm{c}}(\bm{\phi};\lambda_{\mathrm{c}})&=\ I_{\mathrm{\phi}}(\bm{c};\bm{z}_{\mathrm{c}})-\lambda_{\mathrm{c}}I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{c}})\\ \end{split}$ (3.1)
@@
The quantities described in Equation 3.1 are optimized indirectly through a latent bottleneck structure [Tishby2000, Alemi2019, Fischer2020, Moyer2018] combined with adversarial training, as described in the following informal sketch. Note that in the inference model shown in Figure 3(b), the latent variables $\bm{z}_{\mathrm{y}}$ do not depend directly on $\bm{y}$ (and analogously for $\bm{c}$). This is the conditional independence assumption typically used in the Information Bottleneck framework. Intuitively, the encoder is forced to distill the relevant information in $\bm{x}$ that is necessary for reconstructing $\bm{c}$ and $\bm{y}$ into the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. This results in the maximization of the $I_{\mathrm{\phi}}(\bm{c};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{y};\bm{z}_{\mathrm{y}})$ terms in Equation 3.1 during training. The additional requirement of minimizing $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{y}})$ can be satisfied by introducing a Gradient Reversal Layer (GRL) [Ganin2015] at the input of the data-driven decoder component $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$. During optimization, the gradient signal propagated backwards from $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ to the encoder is scaled by $-\lambda$, while the forward pass remains unchanged. Therefore, the GRL can be thought of as a pseudo function $R_{\mathrm{\lambda}}(\bm{z})$ such that $R_{\mathrm{\lambda}}(\bm{z})=\bm{z}$ and $\frac{\mathrm{d}R_{\mathrm{\lambda}}}{\mathrm{d}\bm{z}}=-\lambda\bm{I}$. Positive values of $\lambda$ correspond to adversarial training. Conversely, negative values make the training “collaborative”. The absolute value of $\lambda$ determines the strength of the adversarial or collaborative objective, with larger values corresponding to a stronger regularization effect. By turning the decoder $p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ into an adversary, the GRL penalizes information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that contributes to the reconstruction of $\bm{x}$, biasing the encoder towards representations that are minimally informative about $\bm{x}$.
@@
$\begin{split}\mathcal{L}(\bm{\theta},\bm{\phi};\bm{x},\bm{c},\bm{y})&=\ \mathbb{E}_{q_{\mathrm{\phi}}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})}\left[\log p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})\right]\\ &-D_{\mathrm{KL}}\left(q_{\mathrm{\phi}}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})\ ||\ p(\bm{z}_{\mathrm{x}})p_{\mathrm{\theta}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{c})p_{\mathrm{\theta}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{y})\right)\\ \end{split}$ (3.3)
@@
$\begin{split}\mathcal{L}(\bm{\theta},\bm{\phi},\bm{\psi};\bm{x},\bm{c},\bm{y})&=\mathbb{E}_{q_{\mathrm{\phi}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})}}\left[\alpha_{\mathrm{x}}\log p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})+\alpha_{\mathrm{c}}\log r_{\mathrm{\psi}_{\mathrm{c}}}(\bm{c}|\bm{z}_{\mathrm{c}})+\alpha_{\mathrm{y}}\log r_{\mathrm{\psi}_{\mathrm{y}}}(\bm{y}|\bm{z}_{\mathrm{y}})\right]\\ &-\beta D_{\mathrm{KL}}\left(q_{\mathrm{\phi}}(\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}}|\bm{x})||p(\bm{z}_{\mathrm{x}})p_{\mathrm{\theta}_{\mathrm{c}}}(\bm{z}_{\mathrm{c}}|\bm{c})p_{\mathrm{\theta}_{\mathrm{y}}}(\bm{z}_{\mathrm{y}}|\bm{y})\right)\\ \end{split}$ (3.4)
The loss function described in Equation 3.4 includes additional regularization that can be used to balance the contribution of different terms. These are included for completeness, and are not used in the experiments described in Section 5. A scaling factor $\beta>0$ on the KLD is commonly included in the ELBO as a means of adjusting the strength of the regularization imposed by the KLD term, and to control the capacity of the probabilistic encoder. It is often beneficial to begin training with $\beta=0$ and gradually increase it to $\beta=1$ using an annealing scheme such as the one proposed by [Bowman2016]. Annealing $\beta$ can prevent the model from getting stuck in local minima of the KLD, and the posterior distribution from degenerating to the prior distribution. Conversely, setting $\beta>1$ can promote unsupervised disentanglement [Higgins2016]. Additionally, the log-likelihood function $\log p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, and auxiliary decoders $\log r_{\mathrm{\psi}_{\mathrm{c}}}(\bm{c}|\bm{z}_{\mathrm{c}})$ and $\log r_{\mathrm{\psi}_{\mathrm{y}}}(\bm{y}|\bm{z}_{\mathrm{y}})$ are assigned weights $\alpha_{\mathrm{x}}$, $\alpha_{\mathrm{c}}$ and $\alpha_{\mathrm{y}}$ respectively to allow for balancing the relative strength of these terms. Details of the implementation can be found in Appendix A.
@@ 3.2 Discussion of the approach
The latent bottleneck structure and GRL have several important implications for inference, conditional generation, and uncertainty quantification. These are discussed here, and demonstrated through the case studies presented in Section 5. The main objective of the approach is to ensure that the known physics are properly utilized. Variability in the generative factors will be preferentially captured by the physics-grounded subset of the latent variables. The corresponding posterior distribution will be influenced by the domain and class contributions to the measured response, allowing for domain and class influences to be interpreted through their impact on the physics-grounded latent variables. Therefore, the posterior over physics-grounded latent variables might not necessarily be accurate in the sense of a point estimate of a physical quantity obtained from the posterior being close to the underlying “true value”. This is also in part due to the data-driven component of the decoder, which can yield a constant but not necessarily zero prediction when domain or class influences are not present in the measured response.
@@
The choice of a continuous latent representation for the domain and class variables provides a number of advantages over directly representing the variables themselves, i.e. using the same number and type (e.g. categorical) of latent variables as the domain and class variables. The mapping to a low-dimensional continuous latent space enables the model to deal with high-dimensional domain and class variables, and can improve generalization by promoting the encoding of richer representations of the domain and class [Joy2022]. From an implementation perspective, it is convenient if the decoder inputs are not dependent on the type and dimensionality of the domain and class variables. Furthermore, for discrete and categorical domain and class variables, the latent space enables the model to make a continuous approximation by interpolating over the continuous latent space. Broadly speaking, the continuous latent space allows for more flexibility in the representation of the domain and class variables.
Finally, it is important to note that the proposed architecture is similar to the Domain Invariant Variational Autoencoder (DIVA) proposed by [Ilse2020], but differs in three key aspects. First, DIVA is targeted towards invariant representation learning and domain generalization in a purely data-driven setting. Additionally, DIVA imposes an independence assumption between $\bm{z}_{\mathrm{x}}$, $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. In our proposed approach, the independence assumption is unnecessary due to the adversarial training, which is the third major difference. The lack of an independence assumption facilitates the use of a single probabilistic encoder, potentially reducing the amount of trainable parameters in the model and allowing for more complex and expressive encoder formulations. In our experiments we did not observe a decrease in performance when using a single encoder for all latent variables, compared to using separate encoders for each subset of the latent space, when combined with adversarial training.
3.3 Quantitative assessment of the disentanglement
@@
Quantitatively assessing disentanglement is a challenging problem and several metrics have been proposed [Higgins2016, Chen2018, Kim2018]. This difficulty can be partially attributed to the lack of a consistent definition of disentanglement [Locatello2019]. In practice, the degree of disentanglement achieved by a model is often evaluated based on subjective expectations stemming from domain knowledge [Vowels2019]. Our proposed approach aims to achieve “one-way” disentanglement and is conceptually more akin to techniques that temper the influence of misspecified model components [Carmona2020, Yu2022]. Intuitively, variations in generative factors that are described by the known physics should not affect the data-driven subsets of the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, while variations in generative factors not included in the known physics should still be preferentially captured by the physics-grounded subset of the latent variables. The degree to which this is achieved can be evaluated by comparing the amount of information captured by each subset $\bm{z}_{\mathrm{x}}$, $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ about a specified generative factor. When a subset of the latent variables is informative about a generative factor, it should be possible to train a regressor to predict the value of the generative factor from samples drawn from this subset of the latent variables. Based on this, we propose the following procedure to assess the amount of information about a given generative factor that is encoded in a subset of the latent variables for the trained model:
@@ 4 Previous work
Deep generative models such as VAE describe a mapping between a high-dimensional data manifold, and a low dimensional latent representation. Generative factors in the data are not generally controlled by individual dimensions of the latent variables, nor are they amenable to human interpretation or semantically meaningful. Disentangled representation learning is aimed at learning representations where perturbations of individual dimensions of the latent space correspond to interpretable perturbations of the data [Esmaeili2019]. Approaches for disentangled representation learning can be broadly classified as either unsupervised, where disentanglement is achieved through the use of additional regularization terms on the ELBO, or supervised, semi-supervised, and weakly supervised methods that utilize additional observables or other information. For a comprehensive review of representation learning and of the different approaches, focusing on VAE, the reader is referred to [Bengio2014] and [Tschannen2018] respectively.
@@
Unsupervised disentanglement necessarily relies on inductive bias and implicit supervision [Locatello2019]. A common approach is to adjust the relative importance of the KLD and reconstruction error terms. In the $\beta$-VAE architecture [Higgins2016], the KLD is scaled by a factor $\beta\geq 1$ that determines how much the approximate posterior is penalized for deviating from the prior distribution, limiting the capacity of the latent distribution and forcing independence between the latent variables at the expense of reconstruction quality. Alternative approaches involve weighting the importance of the total correlation term [Watanabe1960], a component of the KLD that quantifies and penalizes dependence between the dimensions of the aggregated posterior distribution (i.e. the posterior distribution marginalized over the entire dataset). To avoid the high computational cost associated with estimating the total correlation, recent works have proposed stochastic approximations based on mini-batches [Chen2018] and adversarial density-ratio estimation [Kim2018]. It has been shown that disentanglement can also be achieved by adjusting the MI between the input and the latent variables, leveraging the information bottleneck theory [Tishby2000]. [Alemi2019] extend this approach to the variational setting, placing a constraint on the MI between the input and latent representation, and through Lagrange relaxation of the resulting constrained optimization problem obtain an objective that encourages the latent representation to act as a sufficient statistic of the input for reconstructing the output. The InfoVAE approach proposed by [Zhao2018] involves scaling specific terms in the ELBO, coupled with an additional term that promotes maximization of the MI between the inputs and latent variables. Several other approaches for unsupervised disentanglement have been proposed in the literature, such as enforcing independence between groups of latent variables [Lopez2018, Esmaeili2019], extending the standard architecture with adversarial components [Larsen2016], additional decoders [Ding2020] and sparsity inducing priors [Tonolini2020].
Several existing works utilize additional observables in a fully supervised [Hadad2018, Sun2022, Debbagh2023], or semi-supervised [Louizos2017] manner, combined with inductive biases in the form of structured models [Siddharth2017], mutual information constraints [Alemi2019] or other regularization methods shared with unsupervised techniques [Achille2017] to promote disentanglement. It has been hypothesized that representation learning approaches can be particularly useful in domain adaptation and transfer learning tasks due to their ability to capture underlying generative factors in data that are shared between tasks [Bengio2014]. The gradient reversal approach utilized in this work was originally proposed to tackle domain adaptation for image classification [Ganin2015, Ganin2016]. Similar approaches have been extended to the setting of multi-view and multi-modal learning [Federici2020, Mondal2023, Aguerri2019, Hwang2020].
Limited guidance is available on incorporating physics-based models in VAE, and particularly on the impact of combining NN components and physics-based models on the interpretability and generalizability of the predictions. In the context of modeling physical systems described by ordinary differential equations, the issue of balancing physics-based and data-driven components is highlighted by [Takeishi2021], and addressed using a regularized training objective based on the posterior predictive check [Gelman2013]. A similar setting is investigated in [Linial2021] and [Yildiz2019], with the latter introducing a regularized objective to ensure consistency of the latent space with the known physics. [Walker2024] present an approach for discovering shared information in multi-modal data where the known physics are utilized to discover modality-dependent features that correlate with some part of the underlying physical process being studied. Recent work has explored the application of adversarial domain adaptation techniques to structural damage identification [Wang2022].

### v3

@@ 3 Proposed Approach
To address the issues presented in Section 2.4, we propose an approach that takes advantage of the domain and class observables to constrain the approximate posterior distribution. Our approach ensures that each subset of the latent variables only encodes information that is relevant to the corresponding modality. This constraint in turn limits the amount of information available to the data-driven components of the decoder, preventing them from correcting every discrepancy between the physics-based model and measurements, and from overriding the known physics. This is achieved by imposing a latent bottleneck structure to the model, combined with an adversarial training objective. A detailed description of the model architecture and derivation of the training objective are provided in Section 3.1, followed by a brief discussion in Section 3.2. A method for quantitatively assessing the information encoded in subsets of the latent variables is presented in Section 3.3.
@@ 3.1 Detailed description of the model
To minimize the reconstruction error, the encoder tends to maximize the information in the posterior distribution over $\bm{z}$ that can be used to predict $\bm{x}$, $\bm{c}$ and $\bm{y}$, subject to the regularization imposed by the prior distribution. For $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, this information includes features from the input signal $\bm{x}$ that are predictive of $\bm{c}$ and $\bm{y}$, but also irrelevant features that are only predictive of $\bm{x}$. These features can include systematic errors stemming from partial knowledge of the physics, and the influence of unknown confounding factors in the measurements. This superfluous information [23], i.e. information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that is not predictive of $\bm{c}$ and $\bm{y}$, can enable the data-driven component of the decoder to override the known physics. Motivated by this observation we aim to simultaneously maximize the information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that is predictive of $\bm{c}$ and $\bm{y}$, while minimizing the information that is predictive of $\bm{x}$. This trade-off can be formalized in terms of the Mutual Information (MI), a measure of the dependence between two random variables [15]. Denoting the MI between $\bm{x}$ and $\bm{z}$ for an encoder parametrized by $\bm{\phi}$ as $I_{\mathrm{\phi}}(\bm{x};\bm{z})$, and introducing the trade-off parameters $\lambda_{\mathrm{c}},\lambda_{\mathrm{y}}$, we define the following relaxed Lagrangian objectives:
[…]
@@
The quantities described in Equation 3.1 are optimized indirectly through a latent bottleneck structure [65, 3, 24, 53] combined with adversarial training, as described in the following informal sketch. Note that in the inference model shown in Figure 3(b), the latent variables $\bm{z}_{\mathrm{y}}$ do not depend directly on $\bm{y}$ (and analogously for $\bm{c}$). This is the conditional independence assumption typically used in the Information Bottleneck framework. Intuitively, the encoder is forced to distill the relevant information in $\bm{x}$ that is necessary for reconstructing $\bm{c}$ and $\bm{y}$ into the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$. This results in the maximization of the $I_{\mathrm{\phi}}(\bm{c};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{y};\bm{z}_{\mathrm{y}})$ terms in Equation 3.1 during training. The additional requirement of minimizing $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{c}})$ and $I_{\mathrm{\phi}}(\bm{x};\bm{z}_{\mathrm{y}})$ can be satisfied by introducing a Gradient Reversal Layer (GRL) [25] at the input of the data-driven decoder component $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$. During optimization, the gradient signal propagated backwards from $g_{\bm{\theta}}(\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ to the encoder is scaled by $\lambda$, while the forward pass remains unchanged. Therefore, the GRL can be thought of as a pseudo function $R_{\mathrm{\lambda}}(\bm{z})$ such that $R_{\mathrm{\lambda}}(\bm{z})=\bm{z}$ and $\frac{\mathrm{d}R_{\mathrm{\lambda}}}{\mathrm{d}\bm{z}}=-\lambda\bm{I}$. Positive values of $\lambda$ correspond to adversarial training. Conversely, negative values make the training “collaborative”. The absolute value of $\lambda$ determines the strength of the adversarial or collaborative objective, with larger values corresponding to a stronger regularization effect. By turning the decoder $p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$ into an adversary, the GRL penalizes information in $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ that contributes to the reconstruction of $\bm{x}$, biasing the encoder towards representations that are minimally informative about $\bm{x}$.
@@
[…]
@@
[…]
The loss function described in Equation 3.4 includes additional regularization hyperparameters that can be used to balance the contribution of different terms. These are included for completeness, and are not used in the experiments described in Section 5. A scaling factor $\beta>0$ on the KLD is commonly included in the ELBO as a means of adjusting the strength of the regularization imposed by the KLD term, and to control the capacity of the probabilistic encoder. It is often beneficial to begin training with $\beta=0$ and gradually increase it to $\beta=1$ using an annealing scheme such as the one proposed by [8]. Annealing $\beta$ can prevent the model from getting stuck in local minima of the KLD, and the posterior distribution from degenerating to the prior distribution. Conversely, setting $\beta>1$ can promote unsupervised disentanglement [32]. The impact of $\beta$ is extensively discussed in the relevant literature, provided in Section 4. Additionally, the log-likelihood function $\log p_{\mathrm{\theta}}(\bm{x}|\bm{z}_{\mathrm{x}},\bm{z}_{\mathrm{c}},\bm{z}_{\mathrm{y}})$, and auxiliary decoders $\log r_{\mathrm{\psi}_{\mathrm{c}}}(\bm{c}|\bm{z}_{\mathrm{c}})$ and $\log r_{\mathrm{\psi}_{\mathrm{y}}}(\bm{y}|\bm{z}_{\mathrm{y}})$ are assigned weights $\alpha_{\mathrm{x}}$, $\alpha_{\mathrm{c}}$ and $\alpha_{\mathrm{y}}$ respectively to allow for balancing the relative strength of these terms (see e.g. [35, 34, 62]). It is important to note that using values other than unity for $\alpha_{\mathrm{x}},\alpha_{\mathrm{c}},\alpha_{\mathrm{y}}$ and $\beta$ can have a significant impact on the interpretation of the ELBO and the inferred posterior distribution. Details of the implementation can be found in Appendix A.
@@ 3.2 Discussion of the approach
The latent bottleneck structure and GRL have several important implications for inference, conditional generation, and uncertainty quantification. These are discussed here, and demonstrated through the case studies presented in Section 5.
3.2.1 Interpretability
The main objective of the approach is to ensure that the known physics are properly utilized, which we interpret as variability in the generative factors being preferentially captured by the physics-grounded subset of the latent variables. As a result, the posterior distribution of the physics-grounded latent variables will be influenced by the domain and class contributions to the measured response, allowing for domain and class influences to be interpreted in terms of their effect on the physics-grounded latent variables. Therefore, the posterior over physics-grounded latent variables might not necessarily be accurate in the sense of a point estimate of a physical quantity obtained from the posterior being close to the underlying “true value”. This is also in part due to the data-driven component of the decoder, which can yield a constant but not necessarily zero prediction when domain or class influences are not present in the measured response.
3.2.2 Conditional generation
@@ 3.2.4 Formulation of the latent space
The choice of a continuous latent representation for the domain and class variables provides a number of advantages over directly representing the variables themselves, i.e. using the same number and type (e.g. categorical) of latent variables as the domain and class variables. The mapping to a low-dimensional continuous latent space enables the model to deal with high-dimensional domain and class variables, and can improve generalization by promoting the encoding of richer representations of the domain and class [35]. From an implementation perspective, it is convenient if the decoder inputs are not dependent on the type and dimensionality of the domain and class variables. Furthermore, for discrete and categorical domain and class variables, the latent space enables the model to make a continuous approximation by interpolating over the continuous latent space. Broadly speaking, the continuous latent space allows for more flexibility in the representation of the domain and class variables. Finally, the lack of an independence assumption facilitates the use of a single probabilistic encoder, potentially reducing the amount of trainable parameters in the model and allowing for more complex and expressive encoder formulations. In our experiments we did not observe a decrease in performance when using a single encoder for all latent variables, compared to using separate encoders for each subset of the latent space, when combined with adversarial training.
3.3 Description of the quantitative assessment approach
@@
Quantitatively assessing disentanglement is a challenging problem and several metrics have been proposed [32, 12, 38]. This difficulty can be partially attributed to the lack of a consistent definition of disentanglement [47]. In practice, the degree of disentanglement achieved by a model is often evaluated based on subjective expectations stemming from domain knowledge [69]. Our proposed approach aims to achieve “one-way” disentanglement and is conceptually more akin to techniques that temper the influence of misspecified model components [11, 76]. Intuitively, variations in generative factors that are described by the known physics should not affect the data-driven subsets of the latent variables $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$, while variations in generative factors not included in the known physics should still be preferentially captured by the physics-grounded subset of the latent variables. The degree to which this is achieved can be evaluated by comparing the amount of information captured by each subset $\bm{z}_{\mathrm{x}}$, $\bm{z}_{\mathrm{c}}$ and $\bm{z}_{\mathrm{y}}$ about a specified generative factor. When a subset of the latent variables is informative about a generative factor, it should be possible to train a regressor to predict the value of the generative factor from samples drawn from this subset of the latent variables. Based on this, we propose the following procedure to assess the amount of information about a given generative factor that is encoded in a subset of the latent variables for the trained model:
@@ 4 Previous work
Deep generative models such as VAE describe a mapping between a high-dimensional data manifold, and a low dimensional latent representation. Generative factors in the data are not generally controlled by individual dimensions of the latent variables, nor are they amenable to human interpretation or semantically meaningful. Disentangled representation learning is aimed at learning representations where perturbations of individual dimensions of the latent space correspond to interpretable perturbations of the data [22]. Approaches for disentangled representation learning can be broadly classified as either unsupervised, where disentanglement is achieved through the use of additional regularization terms on the ELBO, or supervised, semi-supervised, and weakly supervised methods that utilize additional observables or other information. For a comprehensive review of representation learning and of the different approaches, focusing on VAE, the reader is referred to [6] and [67] respectively.
@@
Unsupervised disentanglement necessarily relies on inductive bias and implicit supervision [47]. A common approach is to adjust the relative importance of the KLD and reconstruction error terms. In the $\beta$-VAE architecture [32], the KLD is scaled by a factor $\beta\geq 1$ that determines how much the approximate posterior is penalized for deviating from the prior distribution, limiting the capacity of the latent distribution and encouraging the latent variables to be factorized, at the expense of reconstruction quality [10]. Other approaches involve weighting the importance of the total correlation term [72], a component of the KLD that quantifies and penalizes dependence between the dimensions of the aggregated posterior distribution (i.e. the posterior distribution marginalized over the entire dataset). These approaches avoid the high computational cost associated with estimating the total correlation by utilizing stochastic approximations based on mini-batches [12] and adversarial density-ratio estimation [38]. The InfoVAE approach proposed by [77] involves scaling specific terms in the ELBO, coupled with an additional term that promotes maximization of the MI between the inputs and latent variables. Other approaches for unsupervised disentanglement have been proposed in the literature, such as enforcing independence between and within groups of latent variables [22], extending the standard architecture with adversarial components [44] and additional decoders [20] and using sparsity inducing priors [66]. Several works utilize additional observables in a fully supervised [30, 62, 19], or semi-supervised [49] manner, combined with inductive biases in the form of structured models [54] and penalties on dependence [48] to promote disentanglement or invariance to nuisance factors. It has also been shown [1] that disentanglement is closely related to the information bottleneck theory introduced by [65], and later extended to the variational setting by [3]. Finally, the more general notion of decomposition, that admits disentanglement as a special case, was introduced by [51].
It has been hypothesized that representation learning approaches can be particularly useful in domain adaptation and transfer learning tasks due to their ability to capture underlying generative factors in data that are shared between tasks [6]. The gradient reversal approach utilized in this work was originally proposed to tackle domain adaptation for image classification [25, 26]. Similar approaches have been extended to the setting of multi-view and multi-modal learning [23, 52, 2, 33]. Recent work has also explored the application of adversarial domain adaptation techniques to structural damage identification [71]. Finally, it is important to note that our proposed architecture is similar to the Domain Invariant Variational Autoencoder (DIVA) proposed by [34], from which we adopt part of our terminology. DIVA is targeted towards invariant representation learning and domain generalization in a purely data-driven setting and does not utilize adversarial training, instead explicitly imposing an independence assumption between subsets of the latent variables.
VI offers a balance between accuracy and computational tractability. Combined with the inherent regularization of the Bayesian framework, these properties are particularly advantageous in the modeling of physical systems [27]. As a result, incorporating physical knowledge in VAE has received significant attention, and various physics-informed formulations have been proposed depending on the task of interest. Notably, [70] present an approach for utilizing known physics to discover shared information in multi-modal data. The UQ-VAE [28] combines known governing equations and prior distributions over parameters of interest with paired input-output measurements to achieve computationally efficient uncertainty quantification for systems described by partial differential equations. Formulations of VAE that take advantage of known governing equations have also been proposed for solving forward and inverse problems in stochastic differential equations [78, 61]. In the context of surrogate modeling, [57] introduce the notion of virtual observables as a means of encoding physical knowledge into probabilistic generative models. Despite these advances, the issue of balancing physics-based and data-driven components in VAE has received relatively limited attention. This issue is addressed by [63] for systems described by ordinary differential equations. A similar setting is investigated in [46] and [75], with the latter introducing a regularized objective to ensure consistency of the latent space with the known physics.